Find: (4/9) ÷ (2/3) Find: (4/9) ÷ (2/3) Find: (4/9) ÷ (2/3)

Find:
(4/9) ÷ (2/3)
Find:
(4/9) ÷ (2/3)
Find:
(4/9) ÷ (2/3)

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (4/9) × reciprocal of (2/3)
= (4/9) × (3/2)
= (4 × 3) / (9 × 2)
= (2 × 1) / (3 × 1)
= 2/3

Thanks